# Density-valued time series: Nonparametric density-on-density regression

### v1

@@ 4 Conclusion
Via a series of simulation studies, we show that the proposed method is competitive with several existing methods reported in Kokoszka et al. (2019). Since the true densities are often unobserved in empirical applications, we consider a kernel density estimator to estimate densities. The estimation accuracy of such an estimator depends on the kernel function type and bandwidth parameter selection. For estimating densities, we consider the truncated Gaussian kernel function, and the optimal bandwidth is selected by either Silverman’s ROT. For estimating the nonparametric estimator of the density-on-density regression, the optimal bandwidth parameter is selected by the GCV. With the selected bandwidth, we evaluate and compare the one-step-ahead forecast error between forecast densities constructed based on observed and estimated densities based on a kernel density estimator of the holdout observations using the expanding window approach. For reproducibility, the [Uncaptioned image] code for the forecasting methods and their use in simulation and empirical data analyses are described in https://github.com/hanshang/Bayes_NW.
There are at least three ways the current paper can be further extended: First, a future extension is to develop a bandwidth procedure that jointly selects the optimal bandwidth parameters of the kernel density estimator and Bayes NW estimator. Second, comparing the predictive performance of different methods might not be as fruitful as one might think. It is hard to know when a predictive model works because everything is drowned in the incompressible randomness of new observations. The difference between competing methods can often be seen only in the digits of prediction error, even if they encode substantially different pictures of reality. One may achieve improved forecast accuracy by linearly combining the density forecasts from the nonparametric density-on-density regression and CoDa method. Finally, we proposed a Bayes NW estimator because of its simplicity. Its extension to Bayes local linear estimator can produce improved estimation accuracy.
@@ Acknowledgments
The second author acknowledges financial support from an Australian Research Council Future Fellowship (grant no. FT240100338).
@@
Appendix A Preliminaries
@@
Conditional expectation in $B^{2}(I)$ and model. Let $g$ and $h$ two $\mathcal{B}^{2}(I)$-valued random PDFs; for any $f$ in $\mathcal{B}^{2}(I)$, the conditional expectation $\text{E}(g|h)$ is such that $\langle\text{E}(g|h),f\rangle_{\mathcal{B}}=\text{E}\left(\langle g,f\rangle_{\mathcal{B}}|h\right)$ (see, e.g., Bosq, 2000). Then, $\langle\text{E}(f_{t+1}|f_{t}),g\rangle_{\mathcal{B}}=\text{E}\left(\langle f_{t+1},\,g\rangle_{\mathcal{B}}|f_{t}\right)$ and (1) entails $\langle\text{E}(f_{t+1}|f_{t}),g\rangle_{\mathcal{B}}=\text{E}\left(\langle m(f_{t})\oplus\epsilon,\,g\rangle_{\mathcal{B}}|f_{t}\right)$. With the property of the inner product, $\langle\cdot,\,\cdot\rangle_{\mathcal{B}}$, $\langle\text{E}(f_{t+1}|f_{t}),g\rangle_{\mathcal{B}}=$ $\text{E}\left(\langle m(f_{t}),\,g\rangle_{\mathcal{B}}|f_{t}\right)+\text{E}\left(\langle\epsilon_{t},\,g\rangle_{\mathcal{B}}|f_{t}\right)=\langle m(f_{t}),\,g\rangle_{\mathcal{B}}+\langle\text{E}(\epsilon_{t}|f_{t}),\,g\rangle_{\mathcal{B}}$. According to Model (1), $\text{E}(\epsilon_{t}|f_{t})={\bf 0}_{\mathcal{B}}$ so that, for any $g$ in $\mathcal{B}^{2}(I)$, $\langle\text{E}(f_{t+1}|f_{t}),g\rangle_{\mathcal{B}}=\langle m(f_{t}),g\rangle_{\mathcal{B}}$, which is equivalent to $m(f)=\text{E}\left(f_{t+1}|f_{t}=f\right)$.
Comparing $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|_{\mathcal{B}}^{2}\right)$ and $\text{MISE}(\widehat{f}_{t})$. The next result gives a general situation where $\|f\ominus g\|_{\mathcal{B}}^{2}$ is smaller than $\|f-g\|$ up to a constant.
Proposition 1
Set $\theta>0$; for any $f,g\in\mathcal{B}^{2}(I)$ with $f,g\geq\theta$, $\|f\ominus g\|_{\mathcal{B}}^{2}=O(\|f-g\|^{2})$.
Proof. With the isometric feature of the clr transformation, $\|f\ominus g\|_{\mathcal{B}}^{2}=\int_{I}\{\text{clr}(f)-\text{clr}(g)\}^{2}$, and the definition of clr results in $\|f\ominus g\|_{\mathcal{B}}^{2}=\int_{I}(\log f-\log g)^{2}+\frac{3}{b-a}\left\{\int_{I}(\log f-\log g)\right\}^{2}$. According to the Lipschitz property of the $\log$ function on $[\theta,\,+\infty[$, $|\log f(u)-\log g(u)|\leq\theta^{-1}|f(u)-g(u)|$ as soon as $f(u)\geq\theta$ and $g(u)\geq\theta$ for any $u$; $\|f\ominus g\|_{\mathcal{B}}^{2}\leq M_{\theta}\,\|f-g\|^{2}$ where $M_{\theta}=\theta^{-2}\{1+3/(b-a)\}$. As a by-product, if $f$ is a density in $\mathcal{B}^{2}(I)$ such that $f\geq\theta>0$, then $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|_{\mathcal{B}}^{2}\right)=O\left\{\text{MISE}(\widehat{f}_{t})\right\}$.
Appendix B Proof of Theorem 1
@@
Set $\widehat{K}_{t}=K_{r}\left(h^{-1}\,\|\widehat{f}_{t}\ominus f\|_{\mathcal{B}}\right)$, $\widehat{B}=(N\,\text{E}\widehat{K}_{1})^{-1}\sum^{N}_{t=1}\,\widehat{K}_{t}$, $\widehat{A}=(N\,\text{E}\widehat{K}_{1})^{-1}\odot\bigoplus_{t=1}^{N}\left(\widehat{K}_{t}\odot\widehat{f}_{t+1}\right)$; the proof is based on the decomposition
$\|\widehat{m}_{N}(f)\ominus m(f)\|_{\mathcal{B}}\leq\widehat{B}^{-1}\left\{Q_{1}+Q_{2}\right\}+\widehat{B}^{-1}\|m(f)\|_{\mathcal{B}}\,Q_{3},$ (4)
@@
where $Q_{1}=\|\widehat{A}\ominus\text{E}\widehat{A}\|_{\mathcal{B}}$, $Q_{2}=\|\text{E}\widehat{A}\ominus m(f)\|_{\mathcal{B}}$, and $Q_{3}=|1-\widehat{B}|$. Before going on, let us point out useful features of the Bayes space $\mathcal{B}^{2}(I)$ properties that are systematically used in the proofs. First, the Bayes space of bounded PDFs is a separable Hilbert space; it exists an orthonormal basis $\{e_{j}\}_{j\geq 1}$ such that, for any $g$ in $\mathcal{B}^{2}(I)$, $g=\sum_{j\geq 1}\langle g,e_{j}\rangle_{\mathcal{B}}^{2}$. Second, the standard properties of norms naturally apply in $\mathcal{B}^{2}(I)$; for instance, for any $u_{1}$, $u_{2}$ in $\mathbb{R}$ and any $g_{1}$, $g_{2}$ in $\mathcal{B}^{2}(I)$, $\|u_{1}\odot g_{1}\oplus u_{2}\odot g_{2}\|_{\mathcal{B}}\leq|u_{1}|\|g_{1}\|_{\mathcal{B}}+|u_{2}|\|g_{2}\|_{\mathcal{B}}$. Third, with the assumptions on the kernel function (H(H11)), it exists $C>0$ and $C^{\prime}>0$ such that $C\,\pi_{f}(h_{r})\leq\text{E}\,K_{1}\leq C^{\prime}\pi_{f}(h_{r})$ (resp. $C\,\widehat{\pi}_{f}(h_{r})\leq\text{E}\widehat{K}_{1}\leq C^{\prime}\widehat{\pi}_{f}(h_{r})$), where $\pi_{f}(h_{r})=P\left(\|f_{1}\ominus f\|_{\mathcal{B}}\leq h_{r}\right)$ (resp. $\widehat{\pi}_{f}(h_{r})=P(\|\widehat{f}_{1}\ominus f\|_{\mathcal{B}}\leq h_{r})$). From now on, $C$ and $C^{\prime}$ denote two generic strictly positive constants.
Focus on $Q_{1}$. The separability property of the Bayes space $\mathcal{B}^{2}(I)$ entails the existence of an orthonormal basis $\{e_{j}\}_{j\geq 1}$ such that $\text{E}Q_{1}^{2}=\sum_{j\geq 1}\text{E}\left(\langle\widehat{A}\ominus\text{E}\widehat{A},e_{j}\rangle_{\mathcal{B}}^{2}\right)$. Let $u_{N}$ be a sequence tending to infinity with $N$: $\text{E}Q_{1}^{2}=Q_{1,1}+Q_{1,2}+Q_{1,3}$ with $Q_{1,1}=(N\,\text{E}\widehat{K}_{1})^{-2}\sum_{t=1}^{N}\sum_{j\geq 1}\text{Var}\left(\widehat{K}_{t}\langle\widehat{f}_{t+1},e_{j}\rangle_{\mathcal{B}}\right)$, $Q_{1,2}=(N\,\text{E}\widehat{K}_{1})^{-2}\underset{|s-t|<u_{N}}{\sum_{s=1}^{N}\sum_{t=1}^{N}}\sum_{j\geq 1}\text{cov}_{s,t,j}$, and $Q_{1,3}=(N\,\text{E}\widehat{K}_{1})^{-2}\underset{|s-t|>u_{N}}{\sum_{s=1}^{N}\sum_{t=1}^{N}}\sum_{j\geq 1}\text{cov}_{s,t,j}$, where $\text{cov}_{s,t,j}=\text{Cov}\left(\widehat{K}_{s}\langle\widehat{f}_{s+1},\,e_{j}\rangle_{\mathcal{B}},\,\widehat{K}_{t}\langle\widehat{f}_{t+1},e_{j}\rangle_{\mathcal{B}}\right)$. $Q_{1,1}\leq(N\,\text{E}\widehat{K}_{1})^{-2}\sum_{t=1}^{N}\text{E}\left(\widehat{K}_{t}^{2}\,\|\widehat{f}_{t+1}\|_{\mathcal{B}}^{2}\right)$ and $\|\widehat{f}_{t+1}\|_{\mathcal{B}}\leq\|\widehat{f}_{t+1}\ominus f_{t+1}\|_{\mathcal{B}}+\|f_{t+1}\ominus f\|_{\mathcal{B}}+\|f\|_{\mathcal{B}}$. (H(H5)) and (H(H11)) result in $\|\widehat{f}_{t+1}\|_{\mathcal{B}}=O_{P}(1)$; $Q_{1,1}=O\left(\left\{N\,\widehat{\pi}_{f}(h_{r})\right\}^{-1}\right)$. With the Cauchy-Schwartz inequality and the dependence model, $\sum_{j\geq 1}|\text{cov}_{s,t,j}|\leq 2\,\text{E}\left(\widehat{K}_{1}^{2}\|\widehat{f}_{2}\|_{\mathcal{B}}^{2}\right)$; again, with (H(H5)) and (H(H11)), $Q_{1,2}=O\left(u_{N}\left\{N\,\widehat{\pi}_{f}(h_{r})\right\}^{-1}\right)$. According to the definition of $\rho$-mixing sequence, $\underset{|s-t|>u_{N}}{\sum_{s=1}^{N}\sum_{t=1}^{N}}\sum_{j\geq 1}|\text{cov}_{s,t,j}|\leq\rho(u_{N})\,\text{E}\left(\widehat{K}_{1}^{2}\|\widehat{f}_{2}\|_{\mathcal{B}}^{2}\right)$; then (H(H3)) results in $Q_{1,3}=O\left(u_{N}^{-a}\,\widehat{\pi}_{f}(h_{r})^{-1}\right)$. If we set $u_{N}=N^{1/(a+1)}$, the quantity balancing $Q_{1,2}$ and $Q_{1,3}$, then $Q_{1}=O_{P}\left(\left\{N^{a/(a+1)}\,\widehat{\pi}_{f}(h_{r})\right\}^{-1/2}\right)$.
Focus on $Q_{2}$. Taking into account the dependence model, $Q_{2}=Q_{2,1}+Q_{2,2}+Q_{2,3}$ with $Q_{2,1}=(\text{E}\widehat{K}_{1})^{-1}\text{E}\left(\widehat{K}_{1}\|\widehat{f}_{2}\ominus f_{2}\|_{\mathcal{B}}\right)$, $Q_{2,2}=(\text{E}\widehat{K}_{1})^{-1}\text{E}\left(\widehat{K}_{1}\|m(f_{1})\ominus m(f)\|_{\mathcal{B}}\right)$, $Q_{2,3}=(\text{E}\widehat{K}_{1})^{-1}\left\|\text{E}\left(\widehat{K}_{1}\odot\epsilon_{1}\right)\right\|_{\mathcal{B}}$. Cauchy-Schwartz inequality with (H(H5)) and (H(H11)) entails $\text{E}\left(\widehat{K}_{1}\|\widehat{f}_{2}\ominus f_{2}\|_{\mathcal{B}}\right)=O\left(\widehat{\pi}_{f}(h_{r})^{1/2}\delta_{N}\right)$; use again (H(H11)) to get $Q_{2,1}=O\left(\widehat{\pi}_{f}(h_{r})^{-1/2}\delta_{N}\right)$. With the Lipschitz property of the regression operator $m$ (see (H(H1))), $\|m(f_{1})\ominus m(f)\|_{\mathcal{B}}\leq C\|f_{1}\ominus f\|_{\mathcal{B}}$. Because we only retain $f_{t}$’s such that $K_{t}>0$, $\|m(f_{1})\ominus m(f)\|_{\mathcal{B}}\leq C\,h$ and $Q_{2,2}=O\left(h_{r}\right)$. $\text{E}\left(K_{1}\odot\epsilon_{1}\right)=\text{E}\left\{\text{E}\left(K_{1}\odot\epsilon_{1}|f_{1}\right)\right\}$: $\forall g\in\mathcal{B}^{2}(I)$, $\langle\text{E}\left(K_{1}\odot\epsilon_{1}|f_{1}\right),\,g\rangle_{\mathcal{B}}=K_{1}\,\langle\text{E}\left(\epsilon_{1}|f_{1}\right),\,g\rangle_{\mathcal{B}}=K_{1}\,\langle{\bf 0}_{\mathcal{B}},\,g\rangle_{\mathcal{B}}=0$. Conclusion: for any $g$ in $\mathcal{B}^{2}(I)$, $\langle\text{E}\left(K_{1}\odot\epsilon_{1}|f_{1}\right),\,g\rangle_{\mathcal{B}}=0$, which results in $Q_{2,3}=0$. All these intermediate results in $Q_{2}=O(h_{r})+O\left(\widehat{\pi}_{f}(h_{r})^{-1/2}\delta_{N}\right)$.
Remaining terms and summary. $Q_{3}$ is a particular case of $Q_{1}$ when $\widehat{f}_{t+1}\equiv 1$ and $\widehat{B}=1+(\widehat{B}-1)$ entails $\widehat{B}\leq 1+Q_{3}$ so that $\widehat{B}=1+O_{P}\left(\left\{N^{a/(a+1)}\,\widehat{\pi}_{f}(h_{r})\right\}^{-1/2}\right)$. Then, the decomposition (4) results in

### v2

@@ 4 Conclusion
Through a series of simulation studies, we show that the proposed method is competitive with several existing methods reported in Kokoszka et al. (2019). Since true densities are often unobserved in empirical applications, we consider a kernel density estimator to estimate densities. The estimation accuracy of such an estimator depends on the type of kernel function and the selection of the bandwidth parameter. For estimating densities, we consider the truncated Gaussian kernel function, and the optimal bandwidth is selected by either Silverman’s ROT. For estimating the nonparametric estimator of the density-on-density regression, the optimal bandwidth parameter is selected by the GCV. With the selected bandwidth, we evaluated and compared the one-step-ahead forecast error between the forecast densities constructed based on observed and estimated densities based on a kernel density estimator of the holdout observations using the expanding window approach. For reproducibility, the [Uncaptioned image] code for forecasting methods and their use in simulation and empirical data analyses are described in https://github.com/hanshang/Bayes_NW.
There are at least three ways the current paper can be further extended: First, a future extension is to develop a bandwidth procedure that jointly selects the optimal bandwidth parameters of the kernel density estimator and the Bayes NW estimator. Second, comparing the predictive performance of different methods might not be as fruitful as one might think. It is challenging to determine when a predictive model works because everything is drowned in the incompressible randomness of new observations. The difference between competing methods can often be seen only in the digits of prediction error, even if they encode substantially different pictures of reality. An improved forecast accuracy may be achieved by linearly combining the density forecasts from the nonparametric density-on-density regression and the CoDa method. Finally, we proposed a Bayes NW estimator because of its simplicity. Its extension to the Bayes local linear estimator can produce improved estimation accuracy.
@@ Acknowledgments
The authors thank the editors and two reviewers for their detailed comments, which helped us improve our manuscript. We are grateful for the insightful comments received from the participants at the 4th International Conference on Econometrics and Statistics (EcoSta 2021) and the Functional Data Analysis – A Workshop in Honor of Alois Kneip in 2025. The second author acknowledges financial support from an Australian Research Council Discovery Project (DP230102250) and an Australian Research Council Future Fellowship (FT240100338).
Conflicts of Interest
The authors declare no conflicts of interest.
Data Availability Statement
The data sets that support the findings are publicly available in the French open data portal https://www.data.gouv.fr/fr/ and the Human Mortality Database (2025).
@@
Appendix A Proof of Proposition 1
Proof of (i). Set $\mathcal{B}_{2}(I)^{k}\coloneqq\mathcal{B}_{2}(I)\times\cdots\times\mathcal{B}_{2}(I)$ the Cartesian product of $k$ Bayes Hilbert spaces equipped with norm $\|.\|_{\mathcal{B}^{k}}$ such that $\|(f_{1},\ldots,f_{k})\|_{\mathcal{B}^{k}}\coloneqq\sum_{t=1}^{k}\|f_{t}\|_{\mathcal{B}}$ and let $\mathcal{P}(\mathcal{B}_{2}(I)^{k})$ be the space of probability measures on $\mathcal{B}_{2}(I)^{k}$ with finite second moment. We define the operator $\mathcal{T}$ such that
$\mathcal{T}\,:\,\left\{\begin{array}[]{ccl}\mathcal{P}(\mathcal{B}_{2}(I)^{k})&\rightarrow&\mathcal{P}(\mathcal{B}_{2}(I)^{k})\\ \mu&\mapsto&\mathcal{L}\left((m(f_{1})\oplus\epsilon_{1},\ldots,m(f_{k})\oplus\epsilon_{k})\right),\end{array}\right$.
where $\mathcal{L}\left((m(f_{1})\oplus\epsilon_{1},\ldots,m(f_{k})\oplus\epsilon_{k})\right)$ stands for the law of $(m(f_{1})\oplus\epsilon_{1},\ldots,m(f_{k})\oplus\epsilon_{k})$ and $f\coloneqq(f_{1},\ldots,f_{k})\sim\mu$ (that is, $f$ is drawn by $\mu$). Set $g=(g_{1},\ldots,g_{k})\sim\nu$ with $\nu\in\mathcal{P}(\mathcal{B}_{2}(I)^{k})$; according to the definition of $\|.\|_{\mathcal{B}^{k}}$ and the operators $\oplus$ and $\ominus$,
$\text{E}\|(m(f_{1})\oplus\epsilon_{1},\ldots,m(f_{k})\oplus\epsilon_{k})$ $\ominus_{\mathcal{B}^{k}}(m(g_{1})\oplus\epsilon_{1},\ldots,m(g_{k})\oplus\epsilon_{k})\|_{\mathcal{B}^{k}}$
$=\sum_{t=1}^{k}\text{E}\|(m(f_{t})\oplus\epsilon_{t})\ominus(m(g_{1})\oplus\epsilon_{t})\|_{\mathcal{B}}$
$=\sum_{t=1}^{k}\text{E}\|m(f_{t})\ominus m(g_{t})\|_{\mathcal{B}},$
where $\ominus_{\mathcal{B}^{k}}$ is the perturbation-subtraction operator in the cartesian product $\mathcal{B}(I)^{k}$ such that, for any $k$-tuples $(f_{1},\ldots,f_{k})$ and $(g_{1},\ldots,g_{k})$, $f\ominus_{\mathcal{B}^{k}}g\coloneqq(f_{1}\ominus g_{1},\ldots,f_{k}\ominus g_{k})$. Thanks to the Lipschitz property (H2), there exists $C\in(0,1)$ with, for any $t$,
$\text{E}\|m(f_{t})\ominus m(g_{t})\|_{\mathcal{B}}\leq C\,\text{E}\|f_{t}\ominus g_{t}\|_{\mathcal{B}}.$
As a consequence, there exists $C\in(0,1)$ such that
$\text{E}\|(m(f_{1})\oplus\epsilon_{1},\ldots,m(f_{k})\oplus\epsilon_{k})$ $\ominus_{\mathcal{B}^{k}}(m(g_{1})\oplus\epsilon_{1},\ldots,m(g_{k})\oplus\epsilon_{k})\|_{\mathcal{B}^{k}}$
$\leq C\,\sum_{t=1}^{k}\text{E}\|f_{t}\ominus g_{t}\|_{\mathcal{B}}$
$\leq C\,\text{E}\|f\ominus_{\mathcal{B}^{k}}g\|_{\mathcal{B}^{k}}$ (6)
For $(\mu,\nu)\in\mathcal{B}_{2}(I)^{k}\times\mathcal{B}_{2}(I)^{k}$, let $\Pi(\mu,\nu)$ be the set of probability measures on $\mathcal{B}_{2}(I)^{k}\times\mathcal{B}_{2}(I)^{k}$ having $\mu$ and $\nu$ as marginals and set
$W_{2}\left(\mu,\nu\right)\coloneqq\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathcal{B}_{2}(I)^{k}\times\mathcal{B}_{2}(I)^{k}}\|f\ominus_{\mathcal{B}^{k}}g\|_{\mathcal{B}^{k}}\,d\pi(f,g),$
the Wasserstein 2-distance $W_{2}\left(\mu,\nu\right)$ between $\mu$ and $\nu$. With the definition of operator $\mathcal{T}$ and (6),
$W_{2}\left(\mathcal{T}(\mu),\mathcal{T}(\nu)\right)\leq C\,W_{2}\left(\mu,\nu\right),$
The operator $\mathcal{T}$ is a contraction since $C<1$ (thanks to (H(H3))) and with Banach’s fixed point theorem (see, e.g., Khamsi & Kirk, 2011), there exists a unique probability measure $\mu^{*}$ such that $\mu^{*}=\mathcal{T}(\mu^{*})$. This means that for any $k$, the law of $(f_{1},\ldots,f_{k})$ is equal to the law of $(f_{2},\ldots,f_{k+1})$. The same developments can be iterated $H$ times so that, for any $k$ and $H$, the law of $(f_{1},\ldots,f_{k})$ is equal to the law of $(f_{1+H},\ldots,f_{k+H})$. Conclusion: (H(H3)) implies the existence of a stationary sequence of random PDFs $\{f_{t}\}_{t\in\mathbb{Z}}$ that accommodates the nonparametric model $f_{t+1}=m(f_{t})\oplus\epsilon_{t}$.
To end the proof of $(i)$, just remark that the error $\epsilon_{t}$ is a measurable mapping of ($f_{t},f_{t+1})$ since $clr(\epsilon_{t})=clr(f_{t+1})-clr\circ m(f_{t})$; the $\sigma$-algebra generated by $\epsilon_{p},\ldots,\epsilon_{q}$ (i.e., $\sigma(\epsilon_{t}:\,p\leq t\leq q)$) is equal to $\sigma(f_{t}:\,p\leq t\leq q+1)$. Then, for any $H$, the $\rho$-mixing coefficient $\rho_{\epsilon}(H)$ of $\{\epsilon_{t}\}_{t\in\mathbb{Z}}$ is equal to $\rho_{f}(H-1)$, the $\rho$-mixing coefficient of $\{f_{t}\}_{t\in\mathbb{Z}}$. Consequently, $\{\epsilon_{t}\}_{t\in\mathbb{Z}}$ inherits the $\rho$-mixing property as soon as $\{f_{t}\}_{t\in\mathbb{Z}}$ is $\rho$-mixing and satisfies model (1).
Proof of (ii).
  • About the stationarity of $\{\bm{X}_{t}\}_{t\in\mathbb{Z}}$. For measurable sets $B_{1},\ldots,B_{k}\subset\mathcal{I}^{n}$,
    $\text{Prob}\left(\bm{X}_{t_{1}}\in B_{1},\ldots,\bm{X}_{t_{k}}\in B_{k}\right)=\text{E}\left\{\text{Prob}\left(\bm{X}_{t_{1}}\in B_{1},\ldots,\bm{X}_{t_{k}}\in B_{k}|\{f_{t}\}_{t\in\mathbb{Z}}\right)\right\},$
    and under the conditional probabilistic structure assumed in (H(H1)),
    $\text{Prob}\left(\bm{X}_{t_{1}}\in B_{1},\ldots,\bm{X}_{t_{k}}\in B_{k}\right)=\int_{\Omega}\Pi_{j=1}^{k}\left\{\int_{B_{j}}f_{t_{j}}(\omega,x)^{n}dx\right\}\,d\omega.$
    The right hand side only depends on the law of $f_{t_{1}},\ldots,f_{t_{k}}$ and by stationarity of $\{f_{t}\}_{t\in\mathbb{Z}}$; it equals the analogous expression with each $t_{j}$ shifted by any fixed integer $H$:
    $\text{Prob}\left(\bm{X}_{t_{1}}\in B_{1},\ldots,\bm{X}_{t_{k}}\in B_{k}\right)=\text{Prob}\left(\bm{X}_{t_{1}+H}\in B_{1},\ldots,\bm{X}_{t_{k}+H}\in B_{k}\right).$
  • About the $\rho$-mixing property of $\{\bm{X}_{t}\}_{t\in\mathbb{Z}}$. Set $\mathcal{G}_{p}^{q}=\sigma(\bm{X}_{t}:\,p\leq t\leq q)$ and let $U\in\mathcal{L}^{2}(\mathcal{G}_{-\infty}^{0})$ and $V\in\mathcal{L}^{2}(\mathcal{G}_{H}^{\infty})$ be two rrv with $\text{Var}(U)>0$ and $\text{Var}(V)>0$. Using the same notations introduced to describe the dependence structure (see Section 2.2.3) and the conditional probabilistic structure of $\{\bm{X}_{t}\}_{t\in\mathbb{Z}}$ given process $\{f_{t}\}_{t\in\mathbb{Z}}$ (see (H(H1))),
    $\text{E}(UV)$ $=$ $\text{E}\left\{\text{E}(UV|\{f_{t}\}_{t\in\mathbb{Z}})\right\}$
    $=$ $\text{E}\left(U_{f}V_{f}\right),$
    where $U_{f}=\text{E}(U|\{f_{t}\}_{t\in\mathbb{Z}})$ and $V_{f}=\text{E}(V|\{f_{t}\}_{t\in\mathbb{Z}})$ and we have $\text{Cov}(U,V)=\text{Cov}\left(U_{f},V_{f}\right)$. By the projection property, $\text{Var}(U_{f})\leq\text{Var}(U)$ and $\text{Var}(V_{f})\leq\text{Var}(V)$ so that
    $|\text{Corr}(U,V)|$ $=$ $\frac{|\text{Cov}(U_{f},V_{f})|}{\sqrt{\text{Var}(U)\text{Var}(V)}}$
    $\leq$ $\frac{|\text{Cov}(U_{f},V_{f})|}{\sqrt{\text{Var}(U_{f})\text{Var}(V_{f})}}$
    $\leq$ $|\text{Corr}(U_{f},V_{f})|.$
    Because $U_{f}$ (resp. $V_{f}$) is $\mathcal{F}_{-\infty}^{0}$-measurable (resp. $\mathcal{F}_{H}^{\infty}$-measurable), the definition of $\rho_{f}$ yields $|\text{Corr}(U_{f},V_{f})|\leq\rho_{f}(H)$, which results in $|\text{Corr}(U,V)|\leq\rho_{f}(H)$. The $\rho$-mixing coefficient $\rho_{\bm{X}}(H)$ of $\{\bm{X}_{t}\}_{t\in\mathbb{Z}}$ is such that $\rho_{\bm{X}}(H)\leq\rho_{f}(H)$; $\{\bm{X}_{t}\}_{t\in\mathbb{Z}}$ is a $\rho$-mixing process as soon as the PDF-valued process $\{f_{t}\}_{t\in\mathbb{Z}}$ is a $\rho$-mixing process.
Appendix B Proof of main asymptotic results
B.1 Preliminaries
@@
Conditional expectation in $B^{2}(I)$ and model. Let $g$ and $h$ two $\mathcal{B}_{2}(I)$-valued random PDFs; for any $f$ in $\mathcal{B}_{2}(I)$, the conditional expectation $\text{E}(g|h)$ is such that $\langle\text{E}(g|h),f\rangle_{\mathcal{B}}=\text{E}\left(\langle g,f\rangle_{\mathcal{B}}|h\right)$ (see, e.g., Bosq, 2000). Then, $\langle\text{E}(f_{t+1}|f_{t}),g\rangle_{\mathcal{B}}=\text{E}\left(\langle f_{t+1},\,g\rangle_{\mathcal{B}}|f_{t}\right)$ and (1) entails
$\langle\text{E}(f_{t+1}|f_{t}),g\rangle_{\mathcal{B}}=\text{E}\left(\langle m(f_{t})\oplus\epsilon,\,g\rangle_{\mathcal{B}}|f_{t}\right).$
With the property of the inner product, $\langle\cdot,\,\cdot\rangle_{\mathcal{B}}$,
$\langle\text{E}(f_{t+1}|f_{t}),g\rangle_{\mathcal{B}}$ $=$ $\text{E}\left(\langle m(f_{t}),\,g\rangle_{\mathcal{B}}|f_{t}\right)+\text{E}\left(\langle\epsilon_{t},\,g\rangle_{\mathcal{B}}|f_{t}\right)$
$=$ $\langle m(f_{t}),\,g\rangle_{\mathcal{B}}+\langle\text{E}(\epsilon_{t}|f_{t}),\,g\rangle_{\mathcal{B}}.$
According to model (1), $\text{E}(\epsilon_{t}|f_{t})={\bf 0}_{\mathcal{B}}$ so that, for any $g$ in $\mathcal{B}_{2}(I)$, $\langle\text{E}(f_{t+1}|f_{t}),g\rangle_{\mathcal{B}}=\langle m(f_{t}),g\rangle_{\mathcal{B}}$, which is equivalent to $m(f)=\text{E}\left(f_{t+1}|f_{t}=f\right)$.
B.2 Proof of Theorem 1
@@
Set $\widehat{K}_{t}\coloneqq K_{\text{reg}}\left(h_{\text{reg}}^{-1}\,\|\widehat{f}_{t}\ominus f\|_{\mathcal{B}}\right)$, $\widehat{A}\coloneqq(N\,\text{E}\widehat{K}_{1})^{-1}\odot\bigoplus_{t=1}^{N}\left(\widehat{K}_{t}\odot\widehat{f}_{t+1}\right)$, $\widehat{B}\coloneqq(N\,\text{E}\widehat{K}_{1})^{-1}\sum^{N}_{t=1}\,\widehat{K}_{t}$; the proof is based on the decomposition
$\widehat{m}_{N}(f)\ominus m(f)=\widehat{B}^{-1}\odot\left\{\widehat{A}-\text{E}\widehat{A}\right\}\oplus\widehat{B}^{-1}\odot\left\{\text{E}\widehat{A}\ominus m(f)\right\}\oplus\left(\widehat{B}^{-1}-1\right)\odot m(f),$
which results in
$\|\widehat{m}_{N}(f)\ominus m(f)\|_{\mathcal{B}}\leq\widehat{B}^{-1}\left\{\underbrace{\|\widehat{A}\ominus\text{E}\widehat{A}\|_{\mathcal{B}}}_{Q_{1}}+\underbrace{\|\text{E}\widehat{A}\ominus m(f)\|_{\mathcal{B}}}_{Q_{2}}\right\}+\widehat{B}^{-1}\underbrace{|1-\widehat{B}|}_{Q_{3}}.$ (7)
@@
Before going on, let us point out useful features of the Bayes space $\mathcal{B}_{2}(I)$ that are systematically used in the proofs. First, the Bayes space of bounded PDFs is a separable Hilbert space; there exists an orthonormal basis $\{e_{j}\}_{j\geq 1}$ such that, for any $g$ in $\mathcal{B}_{2}(I)$, $g=\sum_{j\geq 1}\langle g,e_{j}\rangle_{\mathcal{B}}^{2}$. Second, the standard properties of the norms naturally apply in $\mathcal{B}_{2}(I)$; for instance, for any $u_{1}$, $u_{2}$ in $\mathbb{R}$ and any $g_{1}$, $g_{2}$ in $\mathcal{B}_{2}(I)$, $\|u_{1}\odot g_{1}\oplus u_{2}\odot g_{2}\|_{\mathcal{B}}\leq|u_{1}|\|g_{1}\|_{\mathcal{B}}+|u_{2}|\|g_{2}\|_{\mathcal{B}}$. Third, with the assumptions on the kernel function (H(H13)), it exists $C>0$ and $C^{\prime}>0$ such that
$C\,1_{[0,1]}\left(h_{\text{reg}}^{-1}\,\|f_{1}\ominus f\|_{\mathcal{B}}\right)\leq K_{1}\leq C^{\prime}\,1_{[0,1]}\left(h_{\text{reg}}^{-1}\,\|f_{1}\ominus f\|_{\mathcal{B}}\right),$